OpenStudy (anonymous):

A biologist catches 46 fish in a pond, tags them, and then releases them back into the pond. On a later visit to the pond, the biologist catches 20 fish, of which 8 are tagged. Which is the best estimate for the number of fish in the pond?

OpenStudy (anonymous):

He catches 46 different fish one day. The next day he catches 20, but are the same fish he caught before. Do you have an idea where to start?

OpenStudy (anonymous):

but 8 are the same fish*****

OpenStudy (anonymous):

I'm not sure what you mean, the answers are either : A.66 B.115 C.368 D.920

OpenStudy (anonymous):

Well, to solve this we can use a method called Proportion in Lowest Terms.

OpenStudy (anonymous):

Okay.... So how do we do this?

OpenStudy (anonymous):

fish tagged and released = a totalfish = t fish caught again with tags = b fish caught again = c \[\frac{ a }{ t } = \frac{ b }{ c }\]

OpenStudy (anonymous):

a = 46 b = 8 c = 20 t = ? Solve for t

OpenStudy (anonymous):

Is it 920 ?

OpenStudy (anonymous):

46/t = 8/20 ==> \[\frac{ 46 }{ t }=\frac{ 2 }{ 5 }\] \[46=\frac{ 2 }{ 5 }t\] \[46(5)= 2t\] \[t=\frac{ 46(5) }{ 2 }\]

OpenStudy (anonymous):

Oh, 115!

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

no problem